Standard labor hours rate for production of a product are given as follows: Details Hours Rate/hour...

  1. Standard labor hours rate for production of a product are given as follows:

Details

Hours

Rate/hour

Total

Skilled workers

5

1.50

7.50

Unskilled

8

0.50

4.00

Semi-skilled

4

0.75

3.00

            

Actual data                                  Rate per hour            Total

                                                                 $                           $

Output units   1,000

Skilled worker 4,500      hrs                2.00                      9000

Unskilled           10,000   hrs                0.45                      4500

Semi-skilled       4,200   hrs                 0.75                     3150

Required: Calculate

  1. Labor cost variance
  2. Labor rate variance
  3. Labor efficiency variance
  4. Labor mixture variance
  5. Labor yield variance
Answer #1

(a) Labor cost variance = (SH * SR) - (AH * AR)

= [(5000 * 1.5) + (8000 * 0.50) + (4000 * 0.75)] - [(4500 * 2) + ( 10000 * 0.45) + (4200 * 0.75)]

= 14,500 - 16,650

= $ 2,150 (U)

(b) Labor Rate variance = (SR - AR) * AH

= [(1.50 - 2) * 4500] + [(0.50 - 0.45) * 10000] + [(0.75 - 0.75) * 4200]

= $ 1,750 (U)

(c) Labor Efficiency variance = (SH - AH) * SR

= [(5000 - 4500) * 1.50] + [(8000 - 10000) * 0.50] + [(4000 - 4200) * 0.75]

= 400 (U)

(d) Labor Mixture variance = (RAH - AH) * SR

= [(5500 - 4500) * 1.50] + [(8800 - 10000) * 0.50] + [(4400 - 4200) * 0.75]

= $ 1,050 (F)

(e) Labor Yield variance = (SH - RAH) * SR

= [(5000 - 5500) * 1.50] + [(8000 - 8800) * 0.50] + [(4000 - 4400) * 0.75]

= $ 1,450 (U)